OpenStudy (anonymous):

Solve for x. x2 + 6x = 0

OpenStudy (blackops2luvr):

solving x(x-6) =0 then x=0 or x-6=0 so, x=0 0r x=6

OpenStudy (anonymous):

Solve for x. 2x2 - 4x = 0

OpenStudy (blackops2luvr):

no medal? xD

OpenStudy (anonymous):

how do i do that

OpenStudy (blackops2luvr):

click best response

OpenStudy (blackops2luvr):

2x2 – 4x = 0 ; Solve? 2X(X-2)=0 X=2 X=0

OpenStudy (anonymous):

there you go Solve for x. x2 - 2x = 0